INAF queue observing, 2016 Mar 03 - Mar 04

Observer: R. Carini, M. Faccini
Telescope Operator: David Gonzalo-Huerta
Support Astronomer: O. Kuhn on call
Instrument: LUCI

Summary

At the sunset the time is closed

Night duration (12\xB0 el twilight) 19:44-5:27 ( 02:44-12:27 UT) = 10 h 00m

Observing time: 8 h (80%)

Weather time loss: * 2h* (20%)

technical time loss * *m ()

Data Summary

Project_RUN Objects Filter ExpTime Notes
Antoniucci DR Tau 200 H+K grating, zJspec, HKspec filters, ima J 30 m completed in LUCI
tellurica HIP30393 200 H+K grating, zJspec, HKspec filters    
Antoniucci VY Tau 200 H+K grating, zJspec, HKspec filters, ima J 30 m completed in LUCI
Prandoni DDT LH_USS1 200 H+K grating, zJspec, HKspec filters 1h 30m COMPLETED
tellurica HIP53735 200 H+K grating, J, zJspec, HKspec filters
deBarros   200H+K grating, zJspec filter 3 h telluric faint, spectra taken in bad condition of the sky
tellurica HIP59861 200H+K grating, zJspec filter    

Details

00:00 last darks

01:58 bias MODS; flat and arc for Prandoni program slit=1.5 arcsec
  • flat zJspec with tte lamp HALO3 luci2.20160304.0006-0010 texp=30 sec
  • zJspec without lamp luci2.20160304.0011-0015 texp=30 sec
  • arc zJspec wth Neon luci2.20160304.0016-0020 texp=60 sec
  • without lamp luci2.20160304.0021-0025 texp=60 sec
  • arc zJspec wth Argon luci2.20160304.0026-0030 texp=5 sec

02:30 at the beginning of the night the sky is too cloudy to observe.
  • we continue the flats
  • HZKspec without lamp luci2.20160304.0031-0035 texp=30 sec
  • HZKspec HALO3 luci2.20160304.0036-0040 texp=30 sec
  • HZKspec without lamp luci2.20160304.0041-0045 texp=20 sec
  • HZKspec with Neon luci2.20160304.0046-0050 texp=20 sec
  • HZKspec without lamp luci2.20160304.0051-0055 texp=10 sec
  • HZKspec with Argon luci2.20160304.0056-0060 texp=10 sec
  • HZKspec without lamp luci2.20160304.0061-0065 texp=20 sec
  • HZKspec with Xenon luci2.20160304.0066-0070 texp=20 sec
  • zJspec without lamp luci2.20160304.0071-0075 texp=5 sec

03:10 the sky is more clear, configuration LUCI2 * the TO has problems with the configuration * 03:58 begin the collimation, but is too cloudy the collimation failed

04:28 successfully collimation, begin with Antoniucci program, target Dr Tau
  • load xz_tau.xml
  • acquisition luci2.0076-0079
  • seeing ~0.7-''0.8'''
  • science luci2.0080-0087 *4:41 load the script for imaging xz_tau_img.xml luci2.0088-0092
  • seeing 1.2-1.5

04:47 load the script fot the telluric HIP30393
  • acquisition luci2.0093-0096
  • science luci2.0097-0100

05:11 slew to the target VY tau
  • acquisition luci2.0101-0104
  • science luci2.0105-0112
  • imaging luci2.0113-0118
  • seeing 0.8''-0.9'''

05:37 slew to Prandoni DDT
  • 05:39 preset failed. the TO gives the corrections
  • load again the script
  • acquisition luci2.0119-0122
  • the image luci.0123 is a spectrum
  • science luci2.0124-0157
  • seeing ~0.5''-0.7''

07:46 slew to the telluric HIP53735
  • 07:47 error during commit, we load again the script
  • 07:49 error while committing the detector focus, we initialize the detector focus
  • 07:52 load the script, it goes on
  • acquisition luci2.0158-0160
  • science luci2.0161-0172 the last 4 images in the filter HKspec have been done with LS=1.0 arcsec, it should be wrong because the science has been take with LS=1.5 arcsec. We modify the script and load again the script for the Filter HKspec. luci2.0173-0176

08:28 slew to deBarros program
  • acquisition luci2.0177-0180
  • science luci2.0181-0186
  • seeing ~0.4'''

09:22 load again the script of deBarros luci2.0187-0192
  • seeing ~0.4'''

09:56 load the script of deBarros luci2.0193-0198
  • seeing ~0.4''

10:29 load the script of deBarros luci2.0199-0204
  • seeing ~0.6''

11:03 load the script of deBarros luci2.0205-0210
  • seeing ~0.75''
  • 11:15 seeing ~0.5''

11:36 load the script of deBarros luci2.0211-0216
  • seeing variable between 0.5'' and 1.0'', with peaks to 1.2''

12:11 telluric star HIP59861
  • preset failed, the sky is suddenly cloudy.
  • load again the script
  • problem with the collimation
  • 12:34 load again thew script
  • 12:47 THE GUIDING DOESN'T WORK, we do the observation in tracking
  • acquisition luci2.0217-0219
  • science luci2.0220-0226
  • the spectra are present but very faint
  • we take other images increasing the time exposure ==luci2.0227-0229==the situation is not improved

13:30 close the dome

-- %USERSIG{RobertaCarini - 2016-03-04}%

Comments

 
Topic revision: r5 - 09 Mar 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback